# Preempting Fermion Sign Problem: Unveiling Quantum Criticality through Nonequilibrium Dynamics

### v1

@@ II.1 Sign problem behaviors
II.2 Verification of the QCP and critical exponents
Several physical quantities are needed to describe the universal scaling behaviors near the QCP. The ferromagnetic structure factor $S_{\rm FM}$ is defined as
$S_{\rm FM}({\bm{k}})\equiv\frac{1}{L^{2d}}\sum_{i,j}\mathrm{e}^{\mathrm{i}\bm{k}\cdot(\bm{r}_{i}-\bm{r}_{j})}\langle{S_{i}^{z}S_{j}^{z}}\rangle,$ (S9)
@@ II.2 Verification of the QCP and critical exponents
$R_{\rm FM}\equiv 1-\frac{S_{\rm FM}({\bm{k}}={\Delta\bm{k}})}{S_{\rm FM}({\bm{k}}=0)},$ (S10)
where $\Delta\bm{k}=\frac{1}{L}\bm{b_{1}}+\frac{1}{L}\bm{b_{2}}$ is the minimum momentum of electrons in a lattice with periodic boundary conditions, and $\bm{b_{1}},\bm{b_{2}}$ are the reciprocal lattice vectors. The fermion correlation function $G_{\rm f}(\bm{k})$ is defined as
@@
In the main text and figures of data, we abbreviate $S_{\rm FM}\equiv S_{\rm FM}(\bm{k}=\bm{0})$ and $G_{\rm f}\equiv G_{\rm f}(\bm{k}=\Delta\bm{k})$, unless otherwise specified.
@@
In the main text, based on the results from the DSM initial state, we fit according to Eq. (S13) and obtain $U_{c}=7.220(37)$ and $1/\nu=1.18(3)$. Here we supplement with results from the FM initial state, as shown in Fig. S2. Using the same procedure (here we take $\tau L^{-z}=0.5$), we fit according to Eq. (S13) and obtain $U_{c}=7.214(44)$ and $1/\nu=1.05(10)$, which are close to the results from the DSM initial state. These results not only confirm the values of the critical point $U_{c}$ and the critical exponent $1/\nu$, but also show that the initial states can be chosen flexibly in our method, which provides a route to achieve reliable values of critical exponents by benchmarking the results with different initial states.
Figure S2: Verification of the quantum criticality with FM initial state at $\tau=0.5L^{z}$. a, Curves of $R_{\rm FM}$ versus $U$ intersect for different $L$. b, Data collapse after rescaling with $1/\nu=1.18(3)$.
Figure S3: Relaxation dynamics at QCP with FM initial state in single-Dirac-fermion Hubbard model. a-b, Curves of $S_{\rm FM}\equiv S_{\rm FM}(\bm{k}=\bm{0})$ versus $\tau$ for different sizes before and after rescaling. c-d, Curves of $G_{\rm f}\equiv G_{\rm f}(\bm{k}=\Delta\bm{k})$ versus $\tau$ before and after rescaling. Data collapse in the relaxation dynamics shows $\eta_{\phi}=0.34(5)$ and $\eta_{\psi}=0.131(20)$.
$S_{\rm FM}(\bm{k}=\bm{0})=L^{-(1+\eta_{\phi})}f_{S}\left(gL^{1/\nu},\tau L^{-z}\right),$ (S14)
We obtain $\eta_{\phi}=0.34(5)$ and $\eta_{\psi}=0.131(20)$ from the data collapse in the relaxation dynamics at $g=0$, as shown in Fig. S3.
@@
This work (from DSM, $\tau=0.3L^{z}$) 7.220(37) 1.18(3) 0.33(2) 0.135(2)
This work (from FM, $\tau=0.5L^{z}$) 7.214(44) 1.05(10) 0.34(5) 0.131(20)
@@
III More details for the spinless $t$-$V$ model
III.1 Sign problem behaviors
@@
Incidentally, Fig. S4 shows that the decay rate of $\braket{\mathrm{sign}}$ does not monotonically change with the system size $L$. In some cases, a larger system size $L$ results in a weaker sign problem. Such phenomena have also been reported in the equilibrium QMC study [60].
III.2 Verification of the QCP and critical exponents
@@
Here, the sublattice indices $A$ and $B$ function as Dirac spinor indices. In the main text and figures of data, we abbreviate $S_{\rm FM}\equiv S_{\rm FM}(\bm{k}=\bm{0})$ and $G_{\rm f}\equiv G_{\rm f}(\bm{k}=\mathrm{K}+\Delta\bm{k})$, where $\mathrm{K}=\left(\pm\frac{4\pi}{3},0\right)$ represents the momentum at the Dirac points.
@@
Fig. S5 illustrates the correlation length ratio $R_{\rm CDW}$ as a function of interaction strength $V$ at $\tau=0.3L^{z}$. We fit the data in Fig. S5a using the scaling form in Eq. S13, determining the intersection point of the curves for different sizes as $V_{c}=1.37(2)$ and the scaling collapse exponent as $\nu=0.79(5)$. These results are close to those obtained by equilibrium methods [57, 58] and the short-time dynamics with ordered CDW initial state discussed in the main text. Fig. S6 shows the relaxation dynamics of $S_{\rm CDW}=S_{\rm CDW}(\bm{k}=0)$ and $G_{\rm f}=G_{\rm f}(\bm{k}=\mathrm{K}+\Delta\bm{k})$ starting from the DSM initial state. Their scaling behavior is described by Eq. (S14) and Eq. (S15). According to the scaling collapse of the evolution of $S_{\rm CDW}$ in Fig. S6a-S6b, one finds $\eta_{\phi}=0.44(2)$. In addition, from the scaling collapse of the evolution of $G_{\rm f}$ in Fig. S6c-S6d, we obtain $\eta_{\psi}=0.072(4)$. The value of $\eta_{\phi}$ is close to that obtained by the equilibrium methods [57, 58] and the short-time dynamics with ordered CDW initial state discussed in the main text; while the value of $\eta_{\psi}$ is close to the FRG result [56], and also consistent with that obtained from the CDW initial state shown in the main text. Consequently, the results of critical properties in the $t$-$V$ model further demonstrate the efficiency and accuracy of our method. More crucially, for the first time, through our method we achieve the reliable result of fermionic anomalous dimension $\eta_{\psi}$ for the QCP in the spinless $t$-$V$ model by unbiased QMC simulation, as shown in Table S2.
@@
IV More details for the $\rm SU(3)$ Hubbard model
IV.1 $\rm SU(3)$ algebra
@@
$[\lambda_{a},\lambda_{b}]=2\mathrm{i}\sum_{c}f_{abc}\lambda_{c},$ (S20)
@@
IV.2 Mean-field analyses
Figure S7: Mean-field results for the $\lambda_{8}$-AFM order. a, For $\phi=0.1\pi,L=30$, a DSM-AFM transition in the mean-field approximation is seen at $U=3.20(1)$. b, Mean-field phase diagram with $L=30$.
@@ IV.2 Mean-field analyses
After applying the mean-field approximation, the Hamiltonian is entirely expressed as a quadratic form of fermionic operators and can be solved for the ground state using exact diagonalization. The problem is then simply reduced to a self-consistent calculation of the order parameter $\left\{m_{n}\right\}$. Fig. S7a shows the variation of the order parameter $m_{8}$ with interaction strength $U$ calculated using the mean-field method. When the interaction is strong, $m_{8}$ starts to increase with $U$, showing clear characteristics of a continuous phase transition. We control different magnetic flux $\phi$ to find the corresponding phase boundary $U$, as shown in the mean-field phase diagram in Fig. S7b.
@@
whose ground state corresponds to the saturated $\lambda_{n}$-AFM ordered state. At half-filling, the ground state energy of $H^{\text{sat}}_{8}(h)$ is $-\frac{2}{\sqrt{3}}hL^{2}$. The energy gap at half-filling, between the ground state of $H^{\text{sat}}_{8}(h)$ and the first excited state of $H^{\text{sat}}_{8}(h)$, is a finite value $\frac{2}{\sqrt{3}}h$. For $H^{\text{sat}}_{n\neq 8}(h)$, the ground state energy is $-hL^{2}$ and the energy gap is 00. In other words, even considering the saturated ordered state, only the $\lambda_{8}$-AFM order can open a gap in Dirac fermions, while other types of $\lambda_{n}$-AFM cannot.
Figure S8: Mean-field ground state energy per site at half-filling of different types of $\lambda_{n}$-AFM. For the mean-field calculation with $\phi=0.1,L=30$, the ground state energy of the $\lambda_{8}$-AFM mean-field Hamiltonian is lower than that of other types of Hamiltonians.
@@
IV.3 Phase boundary and critical exponents
@@
Figure S10: Correlation-length ratios for $\lambda_{1}$-AFM to $\lambda_{7}$-AFM versus $U$ for different sizes with $\lambda_{8}$-AFM initial state at $\tau=0.25L^{z},\phi=0.075\pi$. There is no crossing point in the curves of $R_{\lambda_{n}\text{-AFM}}$ versus $U$.
@@
To determine the anomalous dimensions $\eta_{\phi}$ and $\eta_{\psi}$ of the bosonic field and the fermionic field we study the relaxation dynamics of the structure factor $S_{\lambda_{8}\text{-AFM}}=S_{\lambda_{8}\text{-AFM}}(\bm{k}=0)$ and the fermion correlation $G_{\rm f}=G_{\rm f}(\bm{k}=\mathrm{K}+\Delta\bm{k})$ at the critical point. The curves before and after rescaling are shown in Figs. S11 and S12 and Fig. 3 in the main text for different $\phi$. The critical exponents determined above are summarized in Table S3, from which we find that the anomalous dimensions for different $\phi$ are consistent with each other within error bar, further confirming the universality of the phase transition.

### v2

@@ DISCUSSION
S2.2 Verification of the QCP and critical exponents
Several physical quantities are needed to describe the universal scaling behaviors near the QCP. The spin structure factor $S({\bm{k}})$ with momentum $\bm{k}$ is defined as
$S({\bm{k}})\equiv\frac{1}{L^{2d}}\sum_{i,j}\mathrm{e}^{\mathrm{i}\bm{k}\cdot(\bm{r}_{i}-\bm{r}_{j})}\langle{S_{i}^{z}S_{j}^{z}}\rangle,$ (S11)
@@
$R_{\rm FM}\equiv 1-\frac{S({\bm{k}}={\Delta\bm{k}})}{S({\bm{k}}=0)},$ (S12)
where $\Delta\bm{k}=\frac{1}{L}\bm{b_{1}}+\frac{1}{L}\bm{b_{2}}$ is the minimum momentum of electrons in a lattice with periodic boundary conditions, and $\bm{b_{1}},\bm{b_{2}}$ are the reciprocal lattice vectors.
The fermion correlation function $G_{\rm f}(\bm{k})$ is defined as
@@
We remark that the fermionic correlation $G_{\rm f}(\bm{k})$ is defined in momentum space, as commonly done in related works [80, 58, 56, 50, 81]. Compared to real-space correlations, which often decay in an oscillatory manner and thus exhibit somewhat subtle finite-size effects (e.g., as it may depend on the even-odd parity of the distance or the size), momentum-space correlations are often more straightforward and robust for analysis. Physically, the momentum-space fermionic correlation defined in Eqs. (S13), (S22) and (S34) captures the quasiparticle weight $Z$, which characterizes the discontinuity in the occupation number across the Fermi surface. Specifically, $Z=\lim_{L\to\infty}2\left|G_{s\bar{s}}(\Delta\bm{k})\right|$, where $s$ and $\bar{s}$ denote the two components of the Dirac spinor, corresponding to spin ($\uparrow,\downarrow$) or sublattice ($A,B$) degrees of freedom in different models. In contrast, momentum-space correlations between the same spinor components are trivial, $G_{ss}(k)=1/2$ for all $k$, due to spontaneous breaking of chiral symmetry. For detailed derivations and discussions, see Refs. [80, 50].
In the main text and figures of data, we abbreviate $S_{\rm FM}\equiv S(\bm{k}=\bm{0})$ and $G_{\rm f}\equiv G_{\rm f}(\bm{k}=\Delta\bm{k})$, unless otherwise specified.
@@
In the main text, based on the results from the DSM initial state, we fit according to Eq. (S15) and obtain $U_{c}=7.220(37),1/\nu=1.18(3)$ at $\tau=0.3L^{z}$. Here we supplement data in Figs. S4c-S4d, where $\tau$ is extended to $0.5L^{z}$, giving $U_{c}=7.225(34),1/\nu=1.16(9)$ at $\tau=0.5L^{z}$, which is consistent with the results at $\tau=0.3L^{z}$ within error bars. This demonstrates that the imaginary time we used has already entered the nonequilibrium scaling regime and that the critical point values have converged. On the other hand, we have also considered the FM initial state, as shown in Figs. S4a-S4b, which gives $U_{c}=7.214(44),1/\nu=1.05(10)$ at $\tau=0.5L^{z}$, also close to the results obtained from the DSM initial state. These results not only confirm the values of the critical point $U_{c}$ and the critical exponent $1/\nu$, but also show that the initial states can be chosen flexibly in our method, which provides a route to achieve reliable values of critical exponents by benchmarking the results with different initial states. This flexibility offers more possible options for alleviating the sign problem.
Figure S4: Verification of quantum criticality using different initial states and imaginary times. ab, With the FM initial state and $\tau=0.5L^{z}$, the data collapse yields $U_{c}=7.214(44)$ and $\nu^{-1}=1.05(10)$. cd, With the DSM initial state and $\tau=0.5L^{z}$, the data collapse yields $U_{c}=7.225(34)$ and $\nu^{-1}=1.16(9)$.
The comparison of results from different initial states, as shown in Table S1, provides an excellent bootstrap-style self-consistency check for the accuracy of our method, which we further illustrate below through the determination of the anomalous dimensions.
@@
This work (from DSM, $\tau=0.3L^{z}$) 7.220(37) 1.18(3) 0.36(3) 0.134(3)
This work (from DSM, $\tau=0.5L^{z}$) 7.225(34) 1.16(9) - -
This work (from FM, $\tau=0.5L^{z}$) 7.214(44) 1.05(10) 0.35(3) 0.136(14)
@@
$S(\bm{k}=\bm{0})=L^{-(1+\eta_{\phi})}f_{S}\left(gL^{1/\nu},\tau L^{-z}\right),$ (S16)
We obtain $\eta_{\phi}=0.34(5)$ and $\eta_{\psi}=0.131(20)$ from the data collapse in the relaxation dynamics at $g=0$, as shown in Fig. S5. Note that only the data within the nonequilibrium scaling region are included in the data collapse analysis. Below we present the general principles and technical details for extracting the critical exponents.
Figure S5: Relaxation dynamics at QCP with FM initial state in single-Dirac-fermion Hubbard model. a-b, Curves of $S_{\rm FM}\equiv S(\bm{k}=\bm{0})$ versus $\tau$ for different sizes before and after rescaling. c-d, Curves of $G_{\rm f}\equiv G_{\rm f}(\bm{k}=\Delta\bm{k})$ versus $\tau$ before and after rescaling. Data collapse in the relaxation dynamics shows $\eta_{\phi}=0.34(5)$ and $\eta_{\psi}=0.131(20)$. Only the interval with sufficiently large $\tau L^{-z}$ is used for the data collapse, with the technical details shown in Fig. S6.
As we stated in Sec. S1.5, when $\tau L^{-z}$ is sufficiently large, the system enters the nonequilibrium scaling region, and the critical exponents can be accurately extracted during the nonequilibrium process without dependence on the initial state. In contrast, for too small values of $\tau L^{-z}$, the system remains too close to the initial state and lies outside the nonequilibrium scaling region governed by the critical point, where the scaling relations no longer hold. In practice, we apply this criterion to precisely determine the critical exponents and delineate the non-equilibrium scaling region. During data collapse analysis, we systematically vary the lower bound of $\tau L^{-z}$ , denoted as $\left(\tau L^{-z}\right)_{\text{min}}$, and examine how the fitted critical exponents vary with $\left(\tau L^{-z}\right)_{\text{min}}$, as shown in Figs. S6a and e. As $\left(\tau L^{-z}\right)_{\text{min}}$ for the data collapse increases, the fitted critical exponents gradually converge and, within the resolution of the error bars, no longer change with $\left(\tau L^{-z}\right)_{\text{min}}$. Moreover, the critical exponents obtained from nonequilibrium processes starting from different initial states converge to the same results within the error bars—this is the hallmark of having entered the nonequilibrium scaling region.
Figure S6: Technical details of determining the critical exponents in the single-Dirac-fermion Hubbard model via data collapse. ad, Determination of the critical exponent $\eta_{\phi}$. a, Fitted values of $\eta_{\phi}$ versus the lower bound $\left(\tau L^{-z}\right)_{\text{min}}$ used in the data collapse analysis. The legend “DSM” denotes data collapse performed with the Dirac semi-metal initial state (the lower set of curves in d); the legend “FM” denotes data collapse performed with the ferromagnetic initial state (the upper set of curves in d). b, Reduced $\chi_{\nu}^{2}$ of the data collapse versus $\left(\tau L^{-z}\right)_{\text{min}}$. c, Distribution of the fitted $\eta_{\phi}$ obtained from 1000 resamplings with $\left(\tau L^{-z}\right)_{\text{min}}=0.3$. The results are $\eta_{\phi}=0.36(3)$ with the DSM initial state and $\eta_{\phi}=0.35(3)$ with the FM initial state. d, Scaling collapse of curves of the structure factor $S_{\rm FM}$ versus rescaled $\tau$ at $U_{c}$. eh, Determination of the critical exponent $\eta_{\psi}$. e, Fitted values of $\eta_{\psi}$ versus $\left(\tau L^{-z}\right)_{\text{min}}$. The legend “DSM” denotes data collapse performed with the DSM initial state (the upper set of curves in h); the legend “FM” denotes data collapse performed with the FM initial state (the lower set of curves in h). f, Reduced $\chi_{\nu}^{2}$ of the data collapse versus $\left(\tau L^{-z}\right)_{\text{min}}$. g, Distribution of the fitted $\eta_{\psi}$ obtained from 1000 resamplings with $\left(\tau L^{-z}\right)_{\text{min}}=0.01$ for the DSM initial state and $\left(\tau L^{-z}\right)_{\text{min}}=0.1$ for the FM initial state. The results are $\eta_{\psi}=0.134(3)$ for the DSM initial state and $\eta_{\psi}=0.136(14)$ for the FM initial state. h, Scaling collapse of curves of the fermion correlation $G_{\rm f}$ versus rescaled $\tau$ at $U_{c}$.
Regarding the critical exponent $\eta_{\phi}$ in the single-Dirac-fermion Hubbard model, Fig. S6a shows that $\eta_{\phi}$ converges once $\left(\tau L^{-z}\right)_{\text{min}}>0.25$, and the results obtained from nonequilibrium processes starting from different initial states are consistent. We use the reduced $\chi_{\nu}^{2}$ to assess the quality of the data collapse:
$\chi_{\nu}^{2}=\frac{1}{\nu}\sum_{i=1}^{N}\sum_{L}^{N_{L}}\frac{\left(y_{iL}-\mu_{i}\right)^{2}}{\Delta y_{iL}^{2}}\,.$ (S18)
For the rescaled curves corresponding to different system sizes $L$ ($N_{L}$ curves in total), we perform linear interpolation and then uniformly sample $N=50$ values of $\tau L^{-z}$ to obtain the curve ordinates $y_{iL}$ and their uncertainties $\Delta y_{iL}$. We then compute the weighted mean $\mu_{i}=\sum_{L}w_{iL}\,y_{iL}/\sum_{L}w_{iL}$ with weights $w_{iL}=1/\Delta y_{iL}^{2}$. The degrees of freedom for the reduced chi-square are $\nu=N(N_{L}-1)$. As shown in Fig. S6b, when $\left(\tau L^{-z}\right)_{\text{min}}$ is very small, $\chi_{\nu}^{2}\gg 1$, indicating very poor collapse quality and that the scaling form cannot describe such short $\left(\tau L^{-z}\right)_{\text{min}}$. As $\left(\tau L^{-z}\right)_{\text{min}}$ increases, $\chi_{\nu}^{2}$ decreases. Around $0.2<\left(\tau L^{-z}\right)_{\text{min}}<0.35$, $\chi_{\nu}^{2}$ approaches $1$, where the quality of the data collapse is optimal. Further increasing $\left(\tau L^{-z}\right)_{\text{min}}$ leads to $\chi_{\nu}^{2}<1$, which implies overfitting. This occurs because data points at larger $\tau L^{-z}$ suffer from more severe sign problems and thus have larger errors, exceeding the resolution of the collapse. Taking into account the convergence behavior in Fig. S6a, the consistency between results from different initial states, and the collapse quality shown in Fig. S6b, we finally choose $\left(\tau L^{-z}\right)_{\text{min}}=0.3$ as the lower bound for $\tau L^{-z}$ in the data collapse analysis of the structure factor $S$ for the single-Dirac-fermion Hubbard model. We evaluate the uncertainty of the critical exponent using a resampling technique. As shown in Fig. S6c, we randomly perturb the data of the structure factor $S$ according to the size of the error bars of the original data and then perform the data collapse analysis again to extract $\eta_{\phi}$. Figure S6c presents the histogram of $\eta_{\phi}$ obtained from 1000 resamplings, showing consistent results between the DSM and FM initial states. By fitting the histogram with a Gaussian distribution, we obtain the critical exponent as $\eta_{\phi}=0.36(3)$ for the DSM initial state and $\eta_{\phi}=0.35(3)$ for the FM initial state. Figure S6d displays the rescaled data collapse using these results. The black dashed line marks $\tau L^{-z}=0.3$, with the region to the right included in the scaling analysis, where curves of different system sizes collapse perfectly. The region to the left of the dashed line is outside the scaling regime, and the deviations from scaling can be seen from the degree of non-overlap between curves. Figure S6d also shows that the system has not yet evolved to equilibrium (which typically requires $\tau L^{-z}\sim 2$33). Nevertheless, we are able to determine the accurate ground-state critical exponents from nonequilibrium data. Comparing Figs. S6a and S6d, for the system sizes we studied, the $\tau L^{-z}$ needed for the convergence of $\eta_{\phi}$ is about one order of magnitude smaller than that required for the convergence of the structure factor $S$. At larger system sizes, the nonequilibrium critical region will be even broader.
The same applies to other similar data collapse figures: our data collapse analysis is only performed using data within the defined scaling range, which appears after a microscopic nonuniversal time scale. Data points at very small $\tau L^{-z}$ lie outside this range and are, therefore, excluded from the analysis. However, it should be emphasized that the this is not the specific issue for the short-time scaling. Even for the more popular finite-size scaling, when the lattice size is too small to enter the scaling region, the finite-size scaling can lose its efficacy. The boundary of this scaling range is critical for the practical application of our method, as it dictates the minimum permissible $\tau L^{-z}$ and thus the extent to which the sign problem can be alleviated. For this reason, we deliberately presented the crossover from scaling violation to scaling satisfaction in these figures. This approach provides a direct visualization of the scaling-range boundary, elucidating our rationale for selecting appropriate $\tau L^{-z}$ values and preventing the misconception that $\tau L^{-z}$ can be chosen arbitrarily small. We explicitly mark the scaling range in these figures with dash lines.
From the above analysis procedure, it is clear that determining critical points and critical exponents using nonequilibrium scaling is highly controllable in terms of accuracy. We can assess the accuracy not only by examining the asymptotic convergence and the quality of the data collapse, but more importantly, by the fact that nonequilibrium evolutions from different initial states are governed by the same ground-state critical exponents. The consistency of the critical points and critical exponents obtained with different initial states is smoking gun evidence that our results are sufficiently accurate and self-consistent. This is a significant advantage of the nonequilibrium approach compared with finite-temperature scaling or finite-size scaling in the conventional equilibrium approaches. Even in situations where the sign problem is particularly severe and prevents accurate results within the scaling range, the early-time results from different initial states can still bracket a controlled range for the critical points and critical exponent.
S2.3 Procedure for determining the dynamical exponent $z$
The quantum phase transitions studied in this work fall into the Dirac QCP universality class, for which the dynamical exponent is known to be $z=1$. However, applying this method to systems with an unknown dynamical exponent requires a self-consistent treatment, as both $z$ and the critical point should be determined together. To address this, we introduce a general procedure for extracting $z$ without prior knowledge of the critical point. We then demonstrate the efficacy of this procedure by applying it to the single-Dirac-fermion Hubbard model, providing a detailed analysis of the numerical results.
Figure S7: Determining the critical point $U_{c}$ and the dynamical exponent $z$ for the single-Dirac-fermion Hubbard model. a, Curves of dimensionless ratio $R_{0}(U,\tau)\equiv S_{\text{FM}}(U,2\tau)/S_{\text{FM}}(U,\tau)$ versus $U$ at different $\tau$. The crossing point gives an initial estimate of the critical point $U_{c}=6.92(1)$. b, Relaxation dynamics of the correlation length ratio $R_{\text{FM}}$ at $U=6.92$. c, Data collapse of b yields an initial estimate of the dynamical exponent $z=1.68(7)$. d, By fixing $\tau=\frac{1}{8}L^{1.68}$, the crossing point of $R_{\text{FM}}$ versus $U$ curves gives an improved estimate of the critical point $U_{c}=7.19(1)$. e, Relaxation dynamics of the correlation length ratio $R_{\text{FM}}$ at $U=7.19$. f, Data collapse of e yields an improved estimate of the dynamical exponent $z=1.21(14)$. g, By fixing $\tau=\frac{1}{4}L^{1.21}$, the crossing point of $R_{\text{FM}}$ versus $U$ curves gives an improved estimate of the critical point $U_{c}=7.23(1)$. h, Relaxation dynamics of the correlation length ratio $R_{\text{FM}}$ at $U=7.23$. i, Data collapse of e yields an improved estimate of the dynamical exponent $z=1.07(9)$.
We introduce a new dimensionless ratio:
$R_{0}(U,\tau)\equiv\frac{S_{\text{FM}}(U,2\tau)}{S_{\text{FM}}(U,\tau)},$ (S19)
where $S_{\text{FM}}$ is the zero-momentum structure factor. Suppose we do not know the precise values of $U_{c}$ and $z$ initially; we determine them via the following iterative steps:
  1. Step 1. We first select a relatively large system size $L$, for instance, $L=23$ in Fig. S7a, such that the factor $\tau L^{-z}$ has negligible influence on the scaling of $R_{0}$. According to the scaling form $R_{0}=2^{\frac{1+\eta_{\phi}}{z}}f_{R_{0}}(g\tau^{\frac{1}{\nu z}})$, the crossing point of $R_{0}$ curves for different $\tau$ provides an initial estimate of the critical point, $U_{c}=6.92(1)$, as shown in Fig. S7a. Note that this estimate of $U_{c}$ may not yet be highly accurate, since the finite-size effects involving $\tau L^{-z}$ have only been approximately neglected.
  2. Step 2. At the tentative critical point $U=6.92$ obtained from Step 1, we estimate the dynamical exponent $z$ from the scaling form of the correlation length ratio, $R_{\text{FM}}=f_{R}(\tau L^{-z})$. In Fig. S7b-S7c, by performing data collapse of the relaxation dynamics of $R_{\text{FM}}$, we obtain an initial estimate for the dynamical exponent: $z=1.68(7)$. Again, this initial estimate may still have uncertainties.
  3. Step 3. Next, we fix the value of $\tau L^{-z}$ using the estimate $z=1.68$ from Step 2 and refine the critical point determination using the nonequilibrium method described in the main text. As shown in Fig. S7d, we obtain an improved estimate of the critical point, $U_{c}=7.19(1)$.
  4. Further iterations. Returning to Step 2, we re-estimate $z$ at the updated critical point $U=7.19$. As shown in Fig. S7e-S7f, this yields an improved value $z=1.21(14)$. One can further iterate Steps 2 and 3 to progressively enhance the accuracy of both $U_{c}$ and $z$.
By repeatedly iterating Steps 2 and 3, the estimates for both $U_{c}$ and $z$ converge towards stable values. In Fig. S7, after three rounds of iterations, we obtain final estimates of $U_{c}=7.23(1)$ and $z=1.07(9)$, which are consistent within error bars with the result $U_{c}=7.220(37)$ presented in the main text and the known Dirac QCP value of $z=1$.
This procedure can also be generally applied to other models and different types of QCP, even when both the location of the critical point and the value of $z$ are initially unknown.
S3 More details for the spinless $t$-$V$ model
S3.1 Sign problem behaviors
Incidentally, Fig. S8 shows that the decay rate of $\braket{\mathrm{sign}}$ does not monotonically change with the system size $L$. In some cases, a larger system size $L$ results in a weaker sign problem. Such phenomena have also been reported in the equilibrium projector QMC study [63] and finite-temperature DQMC [52].
@@
S3.2 Verification of the QCP and critical exponents
@@
Here, the sublattice indices $A$ and $B$ function as Dirac spinor indices. In the main text and figures of data, we abbreviate $S_{\rm CDW}\equiv S_{\rm CDW}(\bm{k}=\bm{0})$ and $G_{\rm f}\equiv G_{\rm f}(\bm{k}=\mathrm{K}+\Delta\bm{k})$, where $\mathrm{K}=\left(\pm\frac{4\pi}{3},0\right)$ represents the momentum at the Dirac points.
@@
Fig. S9 illustrates the correlation length ratio $R_{\rm CDW}$ as a function of interaction strength $V$ at $\tau=0.3L^{z}$. We fit the data in Fig. S9a using the scaling form in Eq. S15, determining the intersection point of the curves for different sizes as $V_{c}=1.37(2)$ and the scaling collapse exponent as $\nu=0.79(5)$. These results are close to those obtained by equilibrium methods [60, 61] and the short-time dynamics with ordered CDW initial state discussed in the main text. Fig. S10 shows the relaxation dynamics of $S_{\rm CDW}=S_{\rm CDW}(\bm{k}=0)$ and $G_{\rm f}=G_{\rm f}(\bm{k}=\mathrm{K}+\Delta\bm{k})$ starting from the DSM initial state. Their scaling behavior is described by Eq. (S16) and Eq. (S17). According to the scaling collapse of the evolution of $S_{\rm CDW}$ in Fig. S10a-S10b, one finds $\eta_{\phi}=0.44(2)$. In addition, from the scaling collapse of the evolution of $G_{\rm f}$ in Fig. S10c-S10d, we obtain $\eta_{\psi}=0.072(4)$. The value of $\eta_{\phi}$ is close to that obtained by the equilibrium methods [60, 61] and the short-time dynamics with ordered CDW initial state discussed in the main text; while the value of $\eta_{\psi}$ is close to the FRG result [59], and also consistent with that obtained from the CDW initial state shown in the main text. Consequently, the results of critical properties in the $t$-$V$ model further demonstrate the efficiency and accuracy of our method. More crucially, for the first time, through our method we achieve the reliable result of fermionic anomalous dimension $\eta_{\psi}$ for the QCP in the spinless $t$-$V$ model by unbiased QMC simulation, as shown in Table S2.
@@
S4 More details for the SU(3) Hubbard model
S4.1 SU(3) algebra
@@
Note that only $\lambda_{8}$ is full rank, the other seven are not. That is to say, only the order associated with $\lambda_{8}$ can gap out all flavors of fermions.
@@
$\lambda_{a},\lambda_{b}]=2\mathrm{i}\sum_{c}f_{abc}\lambda_{c},$ (S24)
@@
S4.2 Mean-field analyses
Figure S11: Mean-field results for the $\lambda_{8}$-AFM order. a, For $\phi=0.1\pi,L=42$, a DSM-AFM transition in the mean-field approximation is seen at $U=3.637(5)$. b, Mean-field phase diagram with $L=42$.
@@
After applying the mean-field approximation, the Hamiltonian is entirely expressed as a quadratic form of fermionic operators and can be solved for the ground state using exact diagonalization. The problem is then simply reduced to a self-consistent calculation of the order parameter $\left\{m_{\lambda_{n}\text{-AFM}}\right\}$. Fig. S11a shows the variation of the order parameter $m_{\lambda_{8}\text{-AFM}}$ with interaction strength $U$ calculated using the mean-field method. When the interaction is strong, $m_{\lambda_{8}\text{-AFM}}$ starts to increase with $U$, showing clear characteristics of a continuous phase transition. We control different magnetic flux $\phi$ to find the corresponding phase boundary $U$, as shown in the mean-field phase diagram in Fig. S11b.
@@
whose ground state corresponds to the saturated $\lambda_{n}$-AFM ordered state. At half-filling, the ground state energy of $H^{\text{sat}}_{8}(h)$ is $\frac{2}{\sqrt{3}}hL^{2}$. The energy gap at half-filling, between the ground state of $H^{\text{sat}}_{8}(h)$ and the first excited state of $H^{\text{sat}}_{8}(h)$, is a finite value $\frac{2}{\sqrt{3}}h$. For $H^{\text{sat}}_{n\neq 8}(h)$, the ground state energy is $hL^{2}$ and the energy gap is 0. In other words, even considering the saturated ordered state, only the $\lambda_{8}$-AFM order can open a gap in Dirac fermions, while other types of $\lambda_{n}$-AFM cannot.
Figure S12: Mean-field ground state energy per site at half-filling of different types of $\lambda_{n}$-AFM. For the mean-field calculation with $\phi=0.1,L=42$, the ground state energy of the $\lambda_{8}$-AFM mean-field Hamiltonian is lower than that of other types of Hamiltonians.
@@
S4.3 Phase boundary and critical exponents
@@
Figure S14: Correlation length ratios for $\lambda_{1}$-AFM to $\lambda_{7}$-AFM versus $U$ for different sizes with $\lambda_{8}$-AFM initial state at $\tau=0.25L^{z},\phi=0.075\pi$. There is no crossing point in the curves of $R_{\lambda_{n}\text{-AFM}}$ versus $U$.
@@
To determine the anomalous dimensions $\eta_{\phi}$ and $\eta_{\psi}$ of the bosonic field and the fermionic field we study the relaxation dynamics of the structure factor $S_{\lambda_{8}\text{-AFM}}=S_{\lambda_{8}\text{-AFM}}(\bm{k}=0)$ and the fermion correlation $G_{\rm f}=G_{\rm f}(\bm{k}=\mathrm{K}+\Delta\bm{k})$ at the critical point. The curves before and after rescaling are shown in Figs. S15 and S16 and Fig. 3 in the main text for different $\phi$. The critical exponents determined above are summarized in Table S3, from which we find that the anomalous dimensions for different $\phi$ are consistent with each other within error bar, further confirming the universality of the phase transition.